# NotebookOS: A Notebook Operating System for Interactive Training with On-Demand GPUs

### v1

@@ 3.1. NotebookOS Overview
Jupyter Server. NotebookOS’s Jupyter Server provides the core services, serving APIs and REST endpoints used by Jupyter web apps, such as Jupyter Notebook (jupyter_notebook, ) and JupyterLab (jupyter_lab_collaborate, ).
Figure 3. Architecture of NotebookOS.
@@
Distributed Kernel. NotebookOS’s Distributed Kernel consists of three replicas, which use the Raft SMR protocol to replicate the CPU-memory state of the IPython process (Step  5 and section 3.2.4). NotebookOS designs a lightweight executor election protocol to elect a proper replica with sufficient GPU resources for running GPU tasks (Step  6 and section 3.2.2).
Distributed Data Store. NotebookOS delegates large object storage and replication using a Distributed Data Store (Step  7). Large objects like model parameters are saved to the Distributed Data Store and asynchronously copied to a different/newly-created replica if NotebookOS detects overload in GPUs and decides to switch to a different executor replica or migrates the current replica to a different GPU server (section 3.2.4).
@@ 3.2.1. Distributed Kernel Creation
Figure 4. Process of creating a new kernel within NotebookOS’s Docker Compose and Docker Swarm modes.
As described in section 3.1, NotebookOS provides a customized Kernel Provisioner to integrate NotebookOS directly with Jupyter. When a user creates a new Distributed Kernel, the Jupyter Server will create a new instance of NotebookOS’s Kernel Provisioner, which issues a StartKernel RPC request to the Global Scheduler (see Step  1 in Figure 4).
@@
After registering with their Local Schedulers, the three kernel replica containers establish peer-to-peer (P2P) connections with one another to run the Raft protocol for transparent kernel SMR. Once the Raft cluster has been created, the kernels notify their Local Schedulers, which in turn notify the Global Scheduler. At this point, the Global Scheduler returns from its StartKernel RPC handler, and the kernel is officially created.
If the Global Scheduler cannot identify three candidate servers to host the replicas of the new Distributed Kernel, then it will invoke a pluggable handler that is set based upon the configured scheduling policy. This invocation will initiate a scale-out operation to provision however many additional servers are required, as detailed in section 3.4.2.
@@ 3.2.2. Distributed Kernel Raft Protocol
Figure 5. NotebookOS’s executor election protocol.
The three replicas of a NotebookOS Distributed Kernel connect to one another directly, forming a peer-to-peer (P2P) network. The replicas use Raft (raft_atc14, ) to perform SMR and use an executor replica election protocol for designating an “executor replica” for cell executions. This protocol is designed to be efficient and fault tolerant: progress can occur even when messages between replicas—or from each replica’s respective Local Scheduler—are dropped or delayed. The protocol is performed by the three replicas each time a user executes a notebook cell. The designated executor replica is responsible for executing the user-submitted code while the other replicas remain idle until the next cell execution request.
Figure 5 provides an overview of the executor election protocol. When a user submits a cell request, the client sends an execute_request message to the Jupyter Server. This message, which encodes the target kernel’s unique ID, is then forwarded to the Global Scheduler. The Global Scheduler then broadcasts a copy of this message to each Local Scheduler managing a server where a kernel replica container is running (see Step  1 in Figure 5). The Global Scheduler may convert the execute_request message to a yield_request message. This instructs the recipient kernel replicas receiving yield_request that they should not attempt to lead the election and instead defer execution to another replica. This message conversion occurs when the Global or Local Scheduler knows that there are not enough resources on a particular server for its kernel replica to execute code.
Upon receiving an execute or yield message, a kernel immediately appends a LEAD or YIELD proposal to its Raft log (Step  2). In Step  3, the kernel replicas wait for all Step  2’s proposals to be committed by Raft to the Raft log. Proposal commitment is handled by the Raft protocol. Kernel replicas will simply take note of any YIELD proposals that are committed, as the election will “fail” if all kernel replicas propose YIELD. This case is described in section 3.2.3.
@@
Once the executor replica finishes executing the cell task, it commits a notification to the Raft log to inform its peers that the execution has finished (Step  7). This notification is thereby received in Step  8. Finally, in Step  9, all replicas send a Jupyter execute_reply message to their Local Scheduler, which forwards the message to the Global Scheduler. These messages are aggregated and merged together by the Global Scheduler before being forwarded back to the Jupyter Server and subsequently the user’s client.
3.2.3. Handling “Failed” Executor Elections
@@
Once the new kernel replica has started and has read the persisted state from the Distributed Data Store, the Global Scheduler terminates the original kernel replica. The Global Scheduler then instructs the remaining replicas to reconfigure their Raft cluster to replace the terminated kernel replica’s “peer address” with that of the newly-created replica. Next, the new replica joins its Raft cluster and begins replaying its Raft log in order to restore its state to that of its peers, after which the Raft cluster of the reconfigured kernel replicas becomes operational again. Finally, the Global Scheduler resubmits the the original execute_request to the migrated kernel replica, ensuring that it is the one to execute the user-submitted code while the other replicas yield.
@@
During kernel replica migrations, the Global Scheduler will query the Container Prewarmer to see if the selected target host has any pre-warmed kernel replica containers available. If so, then a pre-warmed kernel replica container will be used. Pre-warmed containers simply contain a pre-initialized Python runtime with commonly used dependencies, thereby eliminating the on-demand container provisioning overhead. Pre-warming containers with varied runtime dependencies is beyond the scope of this work; however, NotebookOS could leverage existing solutions (sock_atc18, ; vhive_asplos21, ) for this purpose.
@@ 3.2.4. Kernel State Replication
Figure 6. Overview of the Raft-based state synchronization protocol used by NotebookOS’s Distributed Kernel.
NotebookOS’s Distributed Kernel requires an efficient mechanism to maintain consistency across the kernel replicas. To achieve this, the kernels leverages Python AST-based code analysis (python_ast, ) to identify runtime state that needs to be replicated and synchronized via the Raft SMR. An overview of this protocol is shown in Figure 6. As shown in Steps  1 and 2, code submitted for execution is first transformed into an AST by the executor replica before being executed in Step  3. The executor replica then analyzes the Python AST to identify variables that must be synchronized with the executor’s peers, such as global variables storing intermediate state that are likely to be referenced later when executing subsequent notebook cells. After the AST analysis, the executor replica proposes the AST and any relevant state to the Raft log in Step  4. The Raft consensus protocol replicates the update to the distributed log across all replicas. Upon commitment of the newly appended log entry in Step  5, the AST is replayed by the standby replicas, and any changes to standby replicas’ global state are applied to the replica-local variables.
@@ 3.3. GPU Management
NotebookOS leverages dynamic GPU binding at training time to optimize resource management by enabling fine-grained, per-training-task allocations (see step 6 in Figure 3). NotebookOS’s approach to dynamic GPU sharing is also designed to be pluggable, allowing integration with alternative approaches such as GaiaGPU (gaia_gpu, ). We implement a simple approach in NotebookOS’s prototype, with the intent of using a full-featured solution in production.
NotebookOS binds all GPUs available on a server to all its hosted kernel replica containers. Each time an execute_request is submitted by a client, the Global Scheduler embeds the device IDs of the GPUs allocated to the target kernel replica within the request metadata. NotebookOS automatically loads model parameters from the host’s main memory onto all allocated GPUs using the PyTorch API on the critical path of execution requests. This process typically only takes up to a couple hundred milliseconds, so it does not impact performance or interactivity severely. When the executor replica finishes executing the user-submitted code, NotebookOS automatically copies the user’s data from the GPUs to the server’s host memory. The executor replica waits for the GPU compute operations to complete and for in-GPU state to be copied to host memory before returning the result to the user.
After responding to the user’s code submission, the executor replica begins replicating any updated notebook state to the standby replicas. Small state objects are directly replicated using NotebookOS’s Raft SMR protocol. Large objects (model parameters copied from GPU VRAM and training datasets) are asynchronously written to the Distributed Data Store. This process occurs entirely outside of the critical path of the user request and does not impact user experience (section 3.2.4).
@@ 3.4.1. NotebookOS’s Default Placement Policy
Third, NotebookOS implements a dynamic, cluster-wide limit on the SR. While the cluster size is a function of the number of GPUs that are used by actively-training kernel replicas (detailed in section 3.4.2), the maximum permissible SR across the entire cluster is dynamically adjusted. This limit is calculated as $\frac{\sum{S}}{\sum{G}\cdot R}$, where $\sum{S}$ denotes the total number of subscribed GPUs for all kernel replicas across all servers, and $\sum{G}$ represents the total GPU count across all servers. If scheduling an additional kernel replica on a server would cause the server’s SR to exceed this limit, the server is rejected in favor of another. The effect of this cluster-wide SR limit is illustrated in the evaluation results discussed in section 5.3.4.
@@ 3.4.2. Scale-Out Operations
Scale-out operations involve provisioning additional servers in a platform-dependent manner, and then waiting for the Local Schedulers that are started on the new servers to connect and register with the Global Scheduler. Scale-out operations occur in one of two scenarios. First, they are triggered in response to a failed attempt to place one or more replicas of a distributed kernel. Such a failure occurs when there are no viable candidate servers across the cluster to serve the kernel replicas. Upon triggering the scale-out operation, the placement of the corresponding kernel replicas is paused. Resources are immediately reserved for the paused kernel replicas on newly provisioned servers, i.e., before the servers are fully added to the NotebookOS server cluster. Once the servers are ready to serve replica containers, NotebookOS resumes placing the replicas on them.
Second, scale-out operations can also be triggered by NotebookOS’s auto-scaling policy. NotebookOS’s auto-scaler runs on a configurable interval, monitoring cluster resource utilization and determining whether servers should be added or removed. To handle bursts of training requests, NotebookOS maintains a scaling buffer of “extra” servers. To decide if additional servers should be added, the auto-scaler first examines the total number of GPUs actively committed ($\sum C$) to kernel replicas (that are actively executing GPU code) on servers across the NotebookOS cluster, where $C$ is the number of actively-utilized GPUs on a server. The expected cluster capacity ($\sum G^{\prime}$) is defined as $\sum G^{\prime}=f\cdot\sum C$, where $f$ is a scalar multiplier that controls how aggressively NotebookOS scales. If the current cluster capacity is smaller than $\sum G^{\prime}$, additional GPU servers are provisioned. Intuitively, if the number of actively-utilized GPUs is low, then NotebookOS’s auto-scaler will not try to add additional servers. If the number of actively-utilized is GPUs is high and many existing servers are at capacity, then NotebookOS’s auto-scaler triggers scale-out. We set $f$ to 1.05, as we have found empirically that this enables an appropriate degree of auto-scaling during testing.
If cluster resource usage is too low, NotebookOS’s auto-scaler attempts to release 1-2 idle servers at a time (where idle servers are those with no active training kernel replicas.) NotebookOS’s auto-scaler determines if scaling-in is appropriate using a similar approach to scaling-out. Specifically, if $\sum G^{\prime}$ is less than the number of currently-provisioned GPU servers within the cluster, then NotebookOS gradually releases servers—1 to 2 at a time—until this condition is no longer met.
@@ 4. NotebookOS Implementation
NotebookOS is platform-agnostic. We developed and evaluated NotebookOS atop Docker Swarm (docker_swarm, ) and Docker Compose (docker_compose, ), but it can also be deployed atop Kubernetes (kubernetes, ). To simplify deployment, experimentation, and reproducibility, we provide a set of Ansible (ansible, ) playbooks and a Helm (helm, ) chart to automate setup, installation, and deployment on these container orchestration platforms. We have implemented NotebookOS in approximately 282k lines of code over roughly two person-years. Refer to appendix A for additional implementation details.
@@ 5.1.1. Baselines
We compare NotebookOS with several representative baselines that we have implemented within the system itself. Specifically, we implement three additional placement policies, each representing a different category of alternative solutions.
Reservation emulates the behavior of current notebook platforms like Sensei and Google Colab. Reservation creates a single long-running kernel container for each user session. The container remains active for the entire duration of the user session. Fixed resources, including GPUs, are exclusively allocated to each of these long-running kernel containers.
@@
NotebookOS (LCP) (large container pool) is an alternative NotebookOS baseline that sacrifices (some) interactivity in favor of reduced resource cost. NotebookOS (LCP) serves as a tool for exploring the trade-off between interactivity and resource efficiency. NotebookOS (LCP) employs a larger pool of pre-warmed containers than NotebookOS’s default configuration, which uses three kernel replicas and a significantly smaller pre-warmed container pool used during replica migrations. When a cell task arrives, NotebookOS selects a warm container [^1] We analyzed the Python dependencies of notebook containers from SenseiTrace and observed substantial overlaps in common dependencies. from the pool to serve the request. After execution, the container is returned to the pool rather than being terminated, as in NotebookOS’s default policy. Addressing security and privacy concerns related to container sharing is out of scope for this paper but can leverage existing research (trenv_sosp24, ; rainbowcake_asplos24, ; pagurus_atc22, ).

### v2

@@ 3.1. NotebookOS Overview
Jupyter Server. The Jupyter Server provides the core services, serving APIs and REST endpoints used by Jupyter web apps like Jupyter Notebook (jupyter_notebook, ) and JupyterLab (jupyter_lab_collaborate, ).
Figure 3. Architecture overview of NotebookOS.: Architecture of NotebookOS.
@@
Distributed Kernel. NotebookOS’s Distributed Kernel consists of three replicas, which use the Raft SMR protocol to replicate the CPU-memory state of the IPython process (Step 5 and section˜3.2.4). NotebookOS designs a lightweight executor election protocol to elect a proper replica with sufficient GPU resources for running GPU tasks (Step 6 and section˜3.2.2). A replication factor of 5 incurs substantially higher memory, storage, and network cost without delivering significant performance benefit, while a replication factor of 2 is unsupported by the Raft protocol.
Distributed Data Store. NotebookOS offloads large object (e.g., model parameters) storage and replication to a pluggable Distributed Data Store (Step 7). These objects are asynchronously replicated when NotebookOS detects GPU overload and switches to a new executor replica or migrates the current one to a different server (section˜3.2.4). NotebookOS supports Redis (redis, ), AWS S3 (aws_s3, ), and HDFS (hdfs_msst10, ; hopsfs_fast17, ; lambdafs_asplos23, ).
@@ 3.2.1. Distributed Kernel Creation
Figure 4. Process of creating a new kernel within NotebookOS’s Docker Compose and Docker Swarm modes.: Process of creating a new kernel within NotebookOS’s Docker Compose and Docker Swarm modes.
As described in section˜3.1, NotebookOS provides a customized Kernel Provisioner to integrate NotebookOS directly with Jupyter. When a user creates a new Distributed Kernel, the Jupyter Server will create a new instance of NotebookOS’s Kernel Provisioner, which issues a StartKernel RPC request to the Global Scheduler (see Step 1 in Figure 4).
@@
After registering with their Local Schedulers, the three kernel replica containers establish peer-to-peer (P2P) connections with one another to run the Raft protocol for transparent kernel SMR. Once the Raft cluster is established, the kernels notify their Local Schedulers, which notify the Global Scheduler. The Global Scheduler then returns from its StartKernel RPC handler, and the kernel is officially created.
If the Global Scheduler cannot identify three candidate servers to host the replicas of the new Distributed Kernel, then it will invoke a pluggable handler that is set based upon the configured scheduling policy. This invocation will initiate a scale-out operation to provision however many additional servers are required, as detailed in section˜3.4.2.
@@ 3.2.2. Distributed Kernel Raft Protocol
Figure 5. NotebookOS’s executor election protocol.: NotebookOS’s executor election protocol.
The three replicas of a NotebookOS Distributed Kernel connect to one another directly, forming a peer-to-peer (P2P) network. The replicas use Raft (raft_atc14, ) to perform SMR. The use of Raft is motivated by the need for high-availability and durable sessions, even for CPU-only notebooks: existing platforms like Google Colab reclaim idle sessions (colab_faq, ), resulting in the loss of computed state. While recent progress on GPU state checkpoint/restore offers efficient mechanisms for GPU state snapshotting (phoenixos_sosp25, ; modal_gpu_snapshot, ), coarse-grained CPU state checkpoint/restore (criu, ) is often unnecessarily heavyweight. NotebookOS simplifies fault tolerance, replica migration, failover, and the prevention of duplicate cell execution by leveraging the strong consistency guarantees provided by Raft. To this end, NotebookOS adopts a hybrid approach: it performs online replication for small CPU state and uses a distributed data store for asynchronous persistence of large objects.
Kernel replicas perform an executor replica election protocol to designate an “executor replica” each time a user executes a notebook cell. This protocol is designed to be efficient and fault tolerant: progress occurs even when messages between replicas—or from each replica’s respective Local Scheduler—are dropped or delayed. The executor replica is responsible for executing user-submitted code while the other replicas remain idle until the next cell execution request.
Figure 5 provides an overview of the executor election protocol. When a user submits a cell request, the client sends an execute_request message to the Jupyter Server. This message, which encodes the target kernel’s unique ID, is then forwarded to the Global Scheduler. The Global Scheduler then broadcasts a copy of this message to each Local Scheduler managing a server where a kernel replica container is running (see Step 1 in Figure 5). When the Global Scheduler has sufficient resource information, it directly selects the executor replica for a given kernel and bypasses the Raft-based LEAD/YIELD election phase entirely. In this case, it will convert the execute_request message into a yield_request, signaling to the recipient kernel replicas that they should not attempt to participate in the election process and instead defer execution to the designated replica. This message conversion typically occurs when the Global or Local Scheduler determines that a particular server lacks the necessary resources for that server’s kernel replica to execute code.
Upon receiving an execute or yield message, a kernel immediately appends a LEAD or YIELD proposal to its Raft log (Step 2). In Step 3, the kernel replicas wait for all Step 2’s proposals to be committed by Raft to the Raft log. Proposal commitment is handled by the Raft protocol. Kernel replicas will simply take note of any YIELD proposals that are committed, as the election will “fail” if all kernel replicas propose YIELD. This case is described in section˜3.2.3.
@@
Once the executor replica finishes executing the cell task, it commits a notification to the Raft log to inform its peers that the execution has finished (Step 7). This notification is thereby received in Step 8. In Step 9, all replicas send a Jupyter execute_reply message to their Local Scheduler, which forwards the message to the Global Scheduler. These messages are aggregated and merged together by the Global Scheduler before being forwarded back to the Jupyter Server and subsequently the user’s client.
After responding to the user’s code submission, the executor replica begins replicating any updated notebook state to the standby replicas. Small state objects are directly replicated using NotebookOS’s Raft SMR protocol. Large objects (model parameters copied from GPU VRAM and training datasets) are asynchronously written to the Distributed Data Store. This process occurs entirely outside the user request’s critical path, avoiding any impact on user experience (section˜3.2.4).
3.2.3. Handling Failed Executor Elections
@@
Once the new kernel replica has started and has read the persisted state from remote storage, the Global Scheduler terminates the original kernel replica before instructing the remaining replicas to reconfigure their Raft cluster to replace the terminated kernel replica with the newly-created replica. Next, the new replica joins its Raft cluster and begins replaying its Raft log synchronize its state with its peers, after which the Raft cluster becomes operational again. Finally, the Global Scheduler resubmits the execution request to the migrated kernel replica, ensuring that it executes the user-submitted code while the other replicas yield.
@@
During kernel replica migrations, the Global Scheduler will query the Container Prewarmer to see if the selected target host has any pre-warmed containers available. If so, then a pre-warmed container will be used. These containers contain a pre-initialized Python runtime with commonly used dependencies, thereby eliminating the on-demand container provisioning overhead. Pre-warming containers with varied runtime dependencies is beyond the scope of this work; however, NotebookOS could leverage existing checkpoint/restore (sock_atc18, ) or fork-based (vhive_asplos21, ) solutions for this purpose.
@@ 3.2.4. Kernel State Replication
Figure 6. Overview of the Raft-based state synchronization protocol used by NotebookOS’s Distributed Kernel.: Overview of the Raft-based state synchronization protocol used by NotebookOS’s Distributed Kernel.
NotebookOS’s Distributed Kernel requires an efficient mechanism to maintain consistency across the kernel replicas. To achieve this, the kernels leverages Python AST-based code analysis (python_ast, ) to identify runtime state that needs to be replicated and synchronized via the Raft SMR. This procedure can detect and replicate Python-level state as well as state declared in native (C/C++) code, as this state is referenced in the kernel namespace (i.e., the set of user-defined variables stored within the memory of a kernel process). State of external processes or libC cannot be synchronized under the current implementation and is left as future work. An overview of this protocol is shown in Figure 6. As shown in Steps 1 and 2, code submitted for execution is first transformed into an AST by the executor replica before being executed in Step 3. The executor replica then analyzes the Python AST to identify variables that must be synchronized with the executor’s peers, such as global variables storing intermediate state that are likely to be referenced later when executing subsequent notebook cells. After the AST analysis, the executor replica proposes the AST and any relevant state to the Raft log in Step 4. The Raft consensus protocol replicates the update to the distributed log across all replicas. Upon commitment of the newly appended log entry in Step 5, the AST is applied by the standby replicas, and any changes to standby replicas’ global state are applied to the replica-local variables.
@@ 3.3. GPU Management
NotebookOS performs dynamic GPU binding right before a replica begins executing user-submitted code so as to optimize resource management by enabling fine-grained, per-training-task allocations (see step 6 in Figure 3). NotebookOS’s approach to dynamic GPU sharing is also designed to be pluggable, allowing integration with alternative approaches such as GaiaGPU (gaia_gpu, ). We implement a simple approach in NotebookOS’s prototype, with the intent of using a full-featured solution in production.
NotebookOS binds all GPUs available on a server to all its hosted kernel replica containers. Each time a notebook cell is submitted for execution by a client, the Global Scheduler embeds the device IDs of the GPUs allocated to the target kernel replica within the request metadata. NotebookOS automatically loads model parameters from the host’s main memory onto all allocated GPUs using the PyTorch API on the critical path of execution requests. This process typically only takes up to a couple hundred milliseconds, so it does not impact performance or interactivity severely. When the executor replica finishes executing the user-submitted code, NotebookOS automatically copies the user’s data from the GPUs to the server’s host memory. The executor replica returns the result to the user only after the GPU operations finish and GPU state is copied to host memory.
@@ 3.4.1. NotebookOS’s Default Placement Policy
Third, NotebookOS implements a dynamic, cluster-wide limit on the SR. While the cluster size is a function of the number of GPUs used by actively-training kernel replicas (detailed in section˜3.4.2), the maximum permissible SR across the entire cluster is dynamically adjusted. This limit is calculated as $\frac{\sum{S}}{\sum{G}\cdot R}$, where $\sum{S}$ denotes the total number of subscribed GPUs for all kernel replicas across all servers, and $\sum{G}$ represents the total GPU count across all servers. If scheduling an additional kernel replica on a server would cause the server’s SR to exceed this limit, the server is rejected in favor of another. The effect of this cluster-wide SR limit is illustrated in the evaluation results discussed in section˜5.3.4.
@@ 3.4.2. Scale-Out Operations
Scale-out operations involve provisioning additional servers in a platform-dependent manner, and then waiting for the Local Schedulers that are started on the new servers to connect and register with the Global Scheduler. Scale-out operations occur in one of two scenarios. First, they are triggered in response to a failed attempt to place one or more replicas of a distributed kernel. Such a failure occurs when there are no viable candidate servers across the cluster to serve the kernel replicas. Upon triggering the scale-out operation, the placement of the corresponding kernel replicas is paused. Resources are immediately reserved for the paused kernel replicas on newly provisioned servers, i.e., before the servers are fully added to the NotebookOS server cluster. Once the servers are ready, NotebookOS resumes placing the replicas on them.
Second, scale-out operations can also be triggered by NotebookOS’s auto-scaling policy. NotebookOS’s auto-scaler runs on a configurable interval, monitoring cluster resource utilization and determining whether servers should be added or removed. To handle bursts of training requests, NotebookOS maintains a scaling buffer of “extra” servers. To decide if additional servers should be added, the auto-scaler first examines the total number of GPUs actively committed ($\sum C$) to kernel replicas (that are actively executing GPU code) on servers across the NotebookOS cluster, where $C$ is the number of actively-utilized GPUs on a server. The expected cluster capacity ($\sum G^{\prime}$) is defined as $\sum G^{\prime}=f\cdot\sum C$, where $f$ is a scalar multiplier that controls how aggressively NotebookOS scales. If the current cluster capacity is smaller than $\sum G^{\prime}$, additional GPU servers are provisioned. Intuitively, if the number of actively-utilized GPUs is low, then the auto-scaler will not try to add additional servers. If the number of actively-utilized is GPUs is high and many existing servers are at capacity, then NotebookOS’s auto-scaler triggers scale-out. We set $f$ to 1.05, as we have found empirically that this enables an appropriate degree of auto-scaling during testing.
If cluster resource usage is too low, NotebookOS’s auto-scaler attempts to release 1-2 idle servers at a time (where idle servers are those with no active training kernel replicas.) The auto-scaler determines if scaling-in is appropriate using a similar approach to scaling-out: if $\sum G^{\prime}$ is less than the number of currently-provisioned GPU servers within the cluster, then NotebookOS gradually releases servers 1 to 2 at a time until this condition is no longer met.
@@ 4. NotebookOS Implementation
NotebookOS is platform-agnostic. We developed and evaluated NotebookOS atop Docker Swarm (docker_swarm, ) and Docker Compose (docker_compose, ); Kubernetes (kubernetes, ) is also supported. To simplify deployment, experimentation, and reproducibility, we provide a set of Ansible (ansible, ) playbooks to automate setup, installation, and deployment on these container orchestration platforms. We have implemented NotebookOS in approximately 282k lines of code over roughly two person-years. Refer to Appendix C for additional implementation details.
@@ 5.1.1. Baselines
We implement three representative baseline policies within NotebookOS itself, each of which represents a unique category of alternative solutions.
Reservation emulates the behavior of current notebook platforms like Adobe research cluster and Google Colab. Reservation creates one long-running kernel container for each user session that remains active for the entire duration of the user session. Fixed resources, including GPUs, are exclusively allocated to each long-running kernel container.
@@
NotebookOS (LCP) (large container pool) is an alternative NotebookOS baseline that sacrifices (some) interactivity in favor of reduced resource cost. NotebookOS (LCP) serves as a tool for exploring the trade-off between interactivity and resource efficiency. NotebookOS (LCP) employs a larger pool of pre-warmed containers than NotebookOS’s default configuration, which uses three kernel replicas and a significantly smaller pre-warmed container pool used during replica migrations. When a cell task arrives, NotebookOS selects a warm container [^1] We analyzed the Python dependencies of notebook containers from our workload and observed substantial overlaps in common dependencies. from the pool to serve the request. After execution, the container is returned to the pool rather than being terminated, as in NotebookOS’s default policy. Addressing security and privacy concerns related to container sharing is out of scope for this paper but can leverage existing research (trenv_sosp24, ; rainbowcake_asplos24, ; pagurus_atc22, ).